Amy wants to fence in a garden whose length is 2m more than twice its width . What is the greatest width her garden can be, if she has at most 70 m of fencing?
I set it up like this but am not sure if it is right or how to finish it: let w = width
2w + 2 = length
lw < 70
(2w +2) (w) < 70
2w squared + 2w - 70 < 0
Thank you for your help!

Answer by scott8148(5895)

You're slightly astray (just remember, the PERIMETER is 2L+2W). Instead of an inequality, you can find the greatest width by equaling the most fence.

2L+2W=70 ... L+W=35 ... (2W+2)+W=35 ... 3W=33 ... W=11 ... L=24